# Making Parametric Anomaly Detection on Tabular Data Non-Parametric Again

### v1

@@ 3.2. Retrieval methods
(5) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\|\mathbf{e_{z}}-\mathbf{e_{x}}\|$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=\mathbf{e_{x}},$
@@
Second, we consider attention mechanisms to select $\mathcal{H}$. We consider three types of attention inspired by those proposed in (Gorishniy et al., 2023). First, the vanilla attention (later referred to as v-attention) where the score and value function used to select the retrieved samples are defined as
(6) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=W_{Q}(\mathbf{e_{z}})^{\top}W_{K}(\mathbf{e}_{x})$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=W_{V}(\mathbf{e_{x}}).$
@@
Second, we also consider another type of attention module, later referred to as attention-bsim, which involves replacing the score function defined in eq. 6 as follows
(7) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\left\|W_{K}(\mathbf{e_{z}})-W_{K}(\mathbf{e_{x}})\right\|^{2}$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=W_{V}(\mathbf{e_{x}}).$
Third, we consider attention-bsim-bval a modification of the value function in eq. 7 as
(8) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\left\|W_{K}(\mathbf{e_{z}})-W_{K}(\mathbf{e_{x}})\right\|^{2}$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=T(W_{K}(\mathbf{e_{z}})-W_{K}(\mathbf{e_{x}})),$
@@
(9) $\tilde{\mathbf{e}}_{\mathbf{z}}=(1-\lambda)\cdot\mathbf{e_{z}}+\lambda\cdot\frac{1}{k}\sum_{x\in\mathcal{H}}\mathcal{V}(\mathbf{z},\mathbf{x}).$
@@ 3.3. Anomaly score
We rely on a mask bank composed of $m$ $d$-dimensional masks to construct the anomaly score as proposed in (Thimonier et al., 2023a). We apply each mask to each validation sample and reconstruct the masked features to compute the reconstruction error for each mask. Thus, each validation sample is masked and reconstructed $m$ times. The anomaly score is constructed as the average reconstruction error over the $m$ masks. To construct the mask bank, we fix the maximum of features to be masked simultaneously $r$ and and construct $m=\sum_{k=1}^{r}\binom{d}{k}$ masks. Choosing deterministic masks instead of random masks to create the mask bank used for inference is beneficial for two reasons. First, since the model will reconstruct all features at least once, it increases the likelihood of identifying different types of anomalies. Indeed, anomalies that deviate from the normal distribution due to a single feature would only be identified if the corresponding mask hiding this feature would be included. Second, this approach ensures that all samples are masked identically to build the anomaly score. We investigate the impact of constructing a random mask bank instead of a deterministic mask bank in section 5.4.
@@ 4. Experiments
We experiment on an extensive benchmark of tabular datasets following previous work (Shenkar and Wolf, 2022; Thimonier et al., 2023a). The benchmark is comprised of two datasets widely used in the anomaly detection literature, namely Arrhythmia and Thyroid, a second group of datasets, the ”Multi-dimensional point datasets”, obtained from the Outlier Detection DataSets (ODDS) [^3] http://odds.cs.stonybrook.edu/ containing 28 datasets. We also include three real-world datasets from (Han et al., 2022) to replace KDD and KDDRev that were excluded from our benchmark since they have been shown to be flawed (Silva et al., 2020): fraud, campaign, and backdoor. We display each dataset’s characteristics in table 6 in appendix A.1.
Table 2. Comparison of transformer+attention-bsim and transformer+KNN across values of $k$. Here $-1$ stands for $\mathcal{H}=\mathcal{C}$. Some values are N/A either because it is not relevant to compute (e.g. $-1$ for KNN) or when there are not enough samples in the training set for the selected value of $k$.
@@
transformer+KNN
Abalone $42.5\pm$$7.8$ $42.5\pm$$9.5$ $42.3\pm$$7.3$ $42.3\pm$$7.9$ $43.0\pm$$7.9$ $\mathbf{43.3}\pm$$10.3$ N/A
Satellite $65.6\pm$$3.3$ $58.6\pm$$2.9$ $\mathbf{64.7}\pm$$3.8$ $63.3\pm$$2.1$ $63.3\pm$$2.6$ $51.6\pm$$3.3$ N/A
Lympho $88.3\pm$$7.6$ $89.2\pm$$7.9$ $\mathbf{90.0}\pm$$8.2$ $\mathbf{90.0}\pm$$8.2$ N/A N/A N/A
Satimage $89.0\pm$$4.1$ $86.8\pm$$0.4$ $91.7\pm$$3.5$ $\mathbf{92.2}\pm$$2.1$ $91.5\pm$$2.9$ $90.6\pm$$4.0$ N/A
Thyroid $55.5\pm$$4.8$ $55.5\pm$$4.9$ $55.2\pm$$4.2$ $\mathbf{56.1}\pm$$4.7$ $55.8\pm$$3.7$ $54.1\pm$$3.9$ N/A
Cardio $81.0\pm$$4.1$ $65.6\pm$$3.6$ $\mathbf{81.4}\pm$$1.4$ $\mathbf{81.4}\pm$$1.3$ $\mathbf{81.4}\pm$$1.3$ $\mathbf{81.3}\pm$$1.3$ N/A
Ionosphere $88.1\pm$$2.8$ $\mathbf{85.7}\pm$$3.4$ $81.8\pm$$3.0$ $81.6\pm$$2.3$ N/A N/A N/A
mean $70.3$ $69.1$ $\mathbf{72.4}$ $\mathbf{72.4}$ $67.0$ $64.2$ N/A
mean std $4.9$ $4.7$ $4.5$ $4.1$ $3.7$ $4.6$
@@
To evaluate to which extent sample-sample dependencies are relevant for anomaly detection, we compare models that attend to relations between samples to the vanilla transformer model. We compare the different methods discussed the F1-score ($\uparrow$) and AUROC ($\uparrow$) metrics following the literature. For each dataset, we report an average over $20$ runs for $20$ different seeds; we display the detailed results for all five transformer-based methods in tables 8 and 9 in appendix B.1 and report in table 1 the average F1 and AUROC.
We considered three regimes for the transformer dimensions depending on the dataset size. The transformer encoder comprises $2$ or $4$ layers with $4$ attention heads and hidden dimension $e\in\{8,16,32\}$, for smaller to larger datasets. We train the transformer with a mask probability $p_{mask}$ set to $0.25$ or $0.15$ and rely on the LAMB optimizer (You et al., 2020) with $\beta=(0.9,0.999)$ and also included a Lookahead (Zhang et al., 2019) wrapper with slow update rate $\alpha=0.5$ and $k=6$ steps between updates. We also include a dropout regularization with $p=0.1$ for attention weights and hidden layers. We ensure that during training, all samples from a batch are not masked simultaneously so that the retrieval module receives encoded representations of unmasked samples as it will in the inference stage. We considered the same transformer architecture and hyperparameters for the same dataset for each considered approach. For external retrieval modules, we chose for simplicity $\lambda=0.5$ for aggregation as detailed in eq. 9. We study in section 5.2 the effect of varying the value of $\lambda$ on the model’s performance. For the KNN module, we set $k=5$ as the cardinality of $\mathcal{H}$ since KNN-based anomaly detection methods (Ramaswamy et al., 2000) with $k=5$ have shown strong anomaly detection performance (Thimonier et al., 2023a). In contrast, for the attention modules, we set $\mathcal{H}=\mathcal{C}$ and use the attention weights to compute a weighted mean to be aggregated as in equation 9. We further discuss the choice of $k$ in section 5.1. Finally, depending on the dataset, we trained the model until the loss stopped improving for $50$ or $100$ consecutive epochs. We detail in appendix A.2.2 the varying hyperparameters between datasets. Each experiment in the present work can be replicated using the code made available on github [^4] https://github.com/hugothimonier/Making-Parametric-Anomaly-Detection-on-Tabular-Data-Non-Parametric-Again.
@@ Results
As reported in table 1, we observe that not all retrieval modules induce a significant boost in anomaly detection performance. We observe that only the transformer augmented by the attention-bsim module perform significantly better than the vanilla transformer. In deed, augmenting the vanilla transformer with the retrieval module detailed in eq. 7 allows to increase the average F1-score by $4.3$% and AUROC by $1.2$%. This observation contradicts the results obtained for the supervised classification and regression tasks investigated in previous work (Gorishniy et al., 2023) where the module that obtains the best performance is attention-bsim-bval . However, let us mention that the attention-bsim-bval module involved in our work is not identical to the one put forward in (Gorishniy et al., 2023) since it does not involve any label.
@@ 5. Discussion
We randomly selected a subset of the dataset from the benchmark for the ablation study conducted hereafter. We use this subset of datasets in each experiment for computational reasons and to avoid cherry-picking the hyperparameters.
5.1. Choice of $k$
@@
5.2. Choice of $\lambda$

### v2

@@ 3.2. Retrieval methods
(6) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\|\mathbf{h_{z}}-\mathbf{h_{x}}\|$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=\mathbf{h_{x}},$
@@
Second, we consider attention mechanisms to select $\mathcal{H}$. We consider three types of attention inspired by those proposed in (Gorishniy et al., 2023). First, the vanilla attention (later referred to as v-attention), where the score and value function used to select the retrieved samples are defined as
(7) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=W_{Q}(\mathbf{h_{z}})^{\top}W_{K}(\mathbf{h}_{x})$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=W_{V}(\mathbf{h_{x}}).$
@@
Second, we also consider another type of attention module, later referred to as attention-bsim, which involves replacing the score function defined in eq. (7) as follows
(8) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\left\|W_{K}(\mathbf{h_{z}})-W_{K}(\mathbf{h_{x}})\right\|^{2}$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=W_{V}(\mathbf{h_{x}}).$
Third, we consider attention-bsim-bval a modification of the value function in eq. (8) as
(9) $\mathcal{S}(\mathbf{z},\mathbf{x})$ $=-\left\|W_{K}(\mathbf{h_{z}})-W_{K}(\mathbf{h_{x}})\right\|^{2}$
$\mathcal{V}(\mathbf{z},\mathbf{x})$ $=T(W_{K}(\mathbf{h_{z}})-W_{K}(\mathbf{h_{x}})),$
@@
(10) $\tilde{\mathbf{h}}_{\mathbf{z}}=(1-\lambda)\cdot\mathbf{h_{z}}+\lambda\cdot\frac{1}{k}\sum_{x\in\mathcal{H}}\mathcal{V}(\mathbf{z},\mathbf{x}).$
@@ 3.3. Anomaly score
We rely on a mask bank composed of $m$ $d$-dimensional masks to construct the anomaly score. We apply each mask to each validation sample and reconstruct the masked features to compute the reconstruction error for each mask. Thus, each validation sample is masked and reconstructed $m$ times. The anomaly score is constructed as the average reconstruction error over the $m$ masks. To construct the mask bank, we fix the maximum of features to be masked simultaneously $r$ and and construct $m=\sum_{k=1}^{r}\binom{d}{k}$ masks. Choosing deterministic masks instead of random masks to create the mask bank used for inference is beneficial for two reasons. First, since the model will reconstruct all features at least once, it increases the likelihood of identifying different types of anomalies. Indeed, anomalies that deviate from the normal distribution due to a single feature would only be identified if the corresponding mask hiding this feature would be included. Second, this approach ensures that all samples are masked identically to build the anomaly score. We investigate the impact of constructing a random mask bank instead of a deterministic mask bank in section 5.5.
@@
3.4. Training pipeline
Masking
We sample from a Bernoulli distribution with probability $p_{mask}$ whether each of the $d$ features is masked.
$\mbox{mask}=(m_{1},\dots,m_{d}),$
Encoding
For numerical features, we normalize to obtain 00 mean and unit variance, while we use one-hot encoding for categorical features. At this point, each feature $j$ for $j\in[1,2,...,d]$ has an $e_{j}$-dimensional representation, $encoded(\mathbf{x}_{j})\in\mathbb{R}^{e_{j}}$ , where $e_{j}=1$ for numerical features and for categorical features $e_{j}$ corresponds to its cardinality. We then mask each feature according to the sampled mask vector and concatenate each feature representation with the corresponding mask indicator function. Hence, each feature $j$ has an $(e_{j}+1)$-dimensional representation
$((1-m_{j})\cdot encoded(\mathbf{x}_{j}),m_{j})\in\mathbb{R}^{e_{j}+1},$
where $\mathbf{x}_{j}$ is the $j$-th features of sample $\mathbf{x}$.
In-Embedding
We pass each of the features encoded representations of sample $\mathbf{x}$ through learned linear layers $\texttt{Linear}(e_{j}+1,e)$. We also learn $e$-dimensional index and feature-type embeddings as proposed in (Kossen et al., 2021). Both are added to the embedded representation of sample $\mathbf{x}$. The obtained embedded representation is thus of dimension $d\times e$
$h_{\mathbf{x}}=(h_{\mathbf{x}}^{1},h_{\mathbf{x}}^{2},\dots,h_{\mathbf{x}}^{d})\in\mathbb{R}^{d\times e},$
and $h_{\mathbf{x}}^{j}\in\mathbb{R}^{e}$ corresponds to the embedded representation of feature $j$ of sample $\mathbf{x}$.
Transformer Encoder
The embedded representation obtained from the previous step is then passed through a transformer encoder. The output of the transformer $\mathbf{h}_{\mathbf{x}}$ is of dimension $d\times e$.
Out-Embedding
The output of the transformer, $\mathbf{h}_{\mathbf{x}}\in\mathbb{R}^{d\times e}$ is then used to compute an estimation of the original $d$-dimensional vector. To obtain the estimated feature $j$, we take the $j$-th $d$-dimensional representation which is output by the transformer encoder, $h_{\mathbf{x}_{j}}\in\mathbb{R}^{d}$, and pass it through a linear layer $\texttt{Linear}(e,e_{j})$, where $e_{j}$ is $1$ for numerical features and the cardinality for categorical features.
External Retrieval Modules
During training, for a batch $\mathcal{B}$ composed of $b$ samples, for each sample $\mathbf{x}\in\mathcal{B}$, the entire batch serves as the candidates $\mathcal{C}$. In inference, a random subsample of the training set is used as $\mathcal{C}$. Both for training and inference, when possible memory-wise, we use as $\mathcal{B}$ and $\mathcal{C}$ the entire training set.
As input, the retrieval module receives a $\mathbb{R}^{d\times e}$ representation for each sample. Operations described in eq. (6), (7), (8) and (9) are performed on the flattened representations of samples $\mathbf{x}$, $\mathbf{h}^{\mathbf{x}}_{flatten}\in\mathbb{R}^{d\cdot e}$. After selecting $\mathcal{H}\subseteq\mathcal{C}$, each sample is transformed back to its original dimension to allow aggregation as described in eq. (10).
@@ 4. Experiments
We experiment on an extensive benchmark of tabular datasets following previous work (Shenkar and Wolf, 2022; Thimonier et al., 2024). The benchmark is comprised of two datasets widely used in the anomaly detection literature, namely Arrhythmia and Thyroid, a second group of datasets, the ”Multi-dimensional point datasets”, obtained from the Outlier Detection DataSets (ODDS) [^3] http://odds.cs.stonybrook.edu/ containing 28 datasets. We also include three real-world datasets from (Han et al., 2022): fraud, campaign, and backdoor. We display each dataset’s characteristics in table 7 in appendix A.1.
Table 2. Comparison of the F1-score ($\uparrow$) of transformer+attention-bsim across values of $k$. Here $-1$ stands for $\mathcal{H}=\mathcal{C}$. Some values are N/A either because it is not relevant to compute or when there are not enough samples in the training set for the selected value of $k$.
@@
To evaluate to which extent sample-sample dependencies are relevant for anomaly detection, we compare models that attend to relations between samples to the vanilla transformer model. We compare the different methods discussed using the F1-score ($\uparrow$) and AUROC ($\uparrow$) metrics following the literature. For each dataset, we report an average over $20$ runs for $20$ different seeds; we display the detailed results for all five transformer-based methods in tables 8 and 9 in appendix B.1 and report in Table 1 the average F1 and AUROC.
We considered three regimes for the transformer dimensions depending on the dataset size. The transformer encoder comprises $2$ or $4$ layers with $4$ attention heads and hidden dimension $e\in\{8,16,32\}$ for smaller to larger datasets. We train the transformer with a mask probability $p_{mask}$ set to $0.25$ or $0.15$ and rely on the LAMB optimizer (You et al., 2020) with $\beta=(0.9,0.999)$ and also included a Lookahead (Zhang et al., 2019) wrapper with slow update rate $\alpha=0.5$ and $k=6$ steps between updates. We also include a dropout regularization with $p=0.1$ for attention weights and hidden layers. We ensure that during training, all samples from a batch are not masked simultaneously so that the retrieval module receives encoded representations of unmasked samples as it will in the inference stage. We considered the same transformer architecture and hyperparameters for the same dataset for each considered approach. For external retrieval modules, we chose for simplicity $\lambda=0.5$ for aggregation as detailed in eq. (10). We study in section 5.3 the effect of varying the value of $\lambda$ on the model’s performance. For the KNN module, we set $k=5$ as the cardinality of $\mathcal{H}$ since KNN-based anomaly detection methods (Ramaswamy et al., 2000) with $k=5$ have shown strong anomaly detection performance (Shenkar and Wolf, 2022). In contrast, for the attention modules, we set $\mathcal{H}=\mathcal{C}$ and use the attention weights to compute a weighted mean to be aggregated as in eq. (10). We further discuss the choice of $k$ in section 5.2. Finally, depending on the dataset, we trained the model until the loss stopped improving for $50$ or $100$ consecutive epochs. Each experiment in the present work can be replicated using the code made available on github [^4] https://github.com/hugothimonier/Retrieval-Augmented-Deep-Anomaly-Detection-for-Tabular-Data.
@@ Results
As reported in Table 1, we observe that not all retrieval modules induce a significant boost in anomaly detection performance. We observe that only the transformer augmented by the attention-bsim module performs significantly better than the vanilla transformer. Indeed, augmenting the vanilla transformer with the retrieval module detailed in eq. (8) allows to increase the average F1-score by $4.3$% and AUROC by $1.2$%. This result is all the more interesting since it contradicts the results obtained for the supervised classification and regression tasks investigated in previous work (Gorishniy et al., 2023) where the module that obtains the best performance is attention-bsim-bval. However, let us mention that the attention-bsim-bval module involved in our work is not identical to the one put forward in (Gorishniy et al., 2023) as it does not involve any label.
For completeness, we compare the different architectures proposed in the present work to existing methods in the literature. We rely on the experiments conducted in (Thimonier et al., 2024; Shenkar and Wolf, 2022) for the metrics of the competing methods. We display in figure 2 the comparison to existing methods. We compare our methods to recent deep methods, namely GOAD (Bergman and Hoshen, 2020), DROCC (Goyal et al., 2020), NeuTraL-AD (Qiu et al., 2021), the contrastive approach proposed in (Shenkar and Wolf, 2022) and NPT-AD (Thimonier et al., 2024). We also compare to classical non-deep methods such as Isolation Forest (Liu et al., 2008), KNN (Ramaswamy et al., 2000), RRCF (Guha et al., 2016), COPOD (Li et al., 2020) and PIDForest (Gopalan et al., 2019). We refer the reader to (Thimonier et al., 2024) for the detail on the F1-score per dataset for other methods than those shown in Tables 8 and 9 in appendix B.1.
@@ 5. Discussion
5.1. Why combine dependencies?
To account for the fact that the retrieval-augmented transformer outperforms the vanilla transformer, we hypothesize that different types of anomalies require different dependencies to identify them. In this section, we provide a simple example to demonstrate this statement. Consider a simple three dimensional data space, $\mathbf{x}=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$, in which the relation between the features of normal sample are defined as follows,
(12) $\begin{array}[]{cc}x_{2}=&\alpha_{1}+\beta_{1}x_{1}+\varepsilon\\ x_{3}=&\alpha_{2}+\beta_{2}x_{2}^{2}+\varepsilon,\end{array}$
where $\varepsilon$ is some white noise and $(\alpha_{1},\alpha_{2},\beta_{1},\beta_{2})\in\mathbb{R}^{2}$ are scalars.
Let us consider two types of anomalies as shown in Figure 3. First anomalies of type $1$, in which the relations between features are identical to the ones given in eq. (12) but in a different subspace. Now consider type $2$ anomalies, for which the values of the generating feature $x_{1}$ are in the same subspace as normal samples, the relation between $x_{1}$ and $x_{2}$ is the same, but the parameters of the relation between $x_{2}$ and $x_{3}$ differ. Type $1$ (resp. $2$) anomalies are akin to local anomalies (resp. dependency anomalies) discussed in (Han et al., 2022).
Table 3. Share (%) of each class correctly identified ($\uparrow$). Average over $5$ data splits. The Table should be read as follows: On average, the transformer correctly classified $78.5\%$ of type $1$ anomalies as anomalies.
Normal Anomalies Anomalies
(type $1$) (type $2$)
Mask-KNN $93.0\%$ (±$(\pm 0.4)$ $100.0\%$(±$(\pm 0.0)$ $77.3\%$(±$(\pm 4.4)$
Transformer $91.4\%$ (±$(\pm 1.4)$ $78.5\%$(±$(\pm 1.0)$ $100.0\%$(±$(\pm 0.0)$
[…]
To test our hypothesis, we compare the retrieval augmented transformer to the vanilla transformer and Mask-KNN, a reconstruction-based technique introduced in (Thimonier et al., 2024), that relies on KNN imputation to reconstruct masked features. Mask-KNN (resp. the transformer) can be considered approximately equivalent to the retrieval augmented transformer without considering the feature-feature dependencies (resp. the sample-sample dependencies).
In the present framework, models only leveraging inter-feature relations, such as the vanilla transformer, may have limited capacities to identify anomalies if they satisfy the same relations as given in eq. (12) but in a different subspace, i.e., anomalies of type $1$. Similarly, a model that only relies on inter-sample relations, e.g., Mask-KNN, would struggle to correctly identify anomalies of type $2$ as they lie in a subspace close to normal samples.
Synthetic Dataset
We effectively test this hypothesis by constructing a synthetic three-dimensional dataset where the features of normal samples satisfy the relations described in eq. (12). We also construct two types of anomalies where type $1$ anomalies follow the same inter-feature relation as normal samples but with values in a non-overlapping interval as those of the normal class. Type $2$ anomalies are constructed to be close to the normal population but with inter-feature relation differing from eq. (12). This synthetic dataset is displayed in figure 3. The normal population comprises $1000$ samples, and we generate $200$ anomalies for each type. We use half of the normal samples to train models and use the rest merged with the anomalies as the validation set. We compare the capacity of Mask-KNN, the vanilla transformer, and the retrieval-augmented transformer to identify anomalies and display the obtained results in Table 3. We consider the same mask bank, $\{(1,0,0),(0,1,0),(0,0,1)\}$, for all three approaches. We use the same strategy for selecting the decision threshold as detailed in section 4. See appendix B.2 for more details on the experimental setting.
Figure 3. Anomalies of type $1$ ( ) require inter-sample dependencies to be correctly identified with high probability. Anomalies of type $2$ ( ) on the other hand require inter-feature dependencies to be correctly identified.
Results
We observe in Table 3 that the vanilla transformer struggles to detect type $1$ anomalies in comparison with other methods, as they explicitly require inter-sample dependencies to be identified. Nevertheless, it correctly identifies $91.4$% of normal samples on average and perfectly detects type $2$ anomalies. Conversely, Mask-KNN obtains significantly lower performance than competing methods for type $2$ anomalies while correctly identifying type $1$ anomalies and normal samples. Notice how both methods cannot correctly identify anomalies from one of the two anomaly types despite using a perfectly separable dataset. On the contrary, the retrieval augmented transformer can better identify both types of anomalies as it can leverage inter-sample and inter-feature dependencies. This experiment provides information as to why combining dependencies for anomaly detection is relevant: it allows the detection of most anomaly types while other approaches are confined to some anomaly categories.
5.2. Analysis of the effect of the number of helpers $k$ on the performance
Note that we randomly selected a subset of the dataset from the benchmark for the ablation studies conducted hereafter. We use this subset of datasets in each experiment for computational reasons and to avoid cherry-picking the hyperparameters.
@@ 5.2. Analysis of the effect of the number of helpers $k$ on the performance
5.3. Analysis of the effect of the value of $\lambda$